40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Apr 6 02:07:40 2019, Anjali, Update, Frequency noise measurement, Frequency noise measurement of 1 micron source 
    Reply  Wed Apr 10 00:33:09 2019, Anjali, Update, Frequency noise measurement, Frequency noise measurement of 1 micron source Modulation_frequency_100Hz.jpgModulation_frequency_30Hz.jpgModified_setup.JPG
       Reply  Wed Apr 10 16:58:54 2019, rana, Update, IOO, fiber MZ for NPRO freq noise measurements 
          Reply  Wed Apr 10 22:59:22 2019, gautam, Update, IOO, Spooled fiber 
             Reply  Thu Apr 11 09:05:06 2019, Anjali, Update, IOO, Spooled fiber Frequecy_modulation_32_Hz.pdfFrequecy_modulation_140_Hz.pdfFrequecy_modulation_300_Hz.pdfWithout_modulation.pdfNew_fiber_spool.JPG
       Reply  Fri Apr 12 01:22:27 2019, Anjali, Update, Frequency noise measurement, Frequency noise measurement of 1 micron source 
          Reply  Mon Apr 15 22:39:10 2019, gautam, Update, Frequency noise measurement, Alternate setup with PSL pickoff 
             Reply  Wed Apr 17 00:43:38 2019, gautam, Update, Frequency noise measurement, MZ interferometer ---> DAQ 
                Reply  Thu Apr 25 03:32:25 2019, Anjali, Update, Frequency noise measurement, MZ interferometer ---> DAQ Time_domain_output.pdfFrequency_noise.pdfschematic_heterodyne_setup.pngNoise_budget_1_micron_in_Hz_per_rtHz.pdf
                   Reply  Thu Apr 25 10:25:19 2019, gautam, Update, Frequency noise measurement, Homodyne v Heterodyne 
                      Reply  Thu Apr 25 15:47:54 2019, Anjali, Update, Frequency noise measurement, Homodyne v Heterodyne Experimental_setup.JPG
Message ID: 14540     Entry time: Fri Apr 12 01:22:27 2019     In reply to: 14529     Reply to this: 14544
Author: Anjali 
Type: Update 
Category: Frequency noise measurement 
Subject: Frequency noise measurement of 1 micron source 

The alignement was disturbed after the replcement of the beam splitter. We tried to get the alignment back . But we are not succeeded yet in getting good interfernce pattern. This is mainly because of poor mode matching of two beams. We will also try with the spooled fiber.

Quote:

 

  • The experimental setup is then modified as shown in attachment #3. The thick beam spliiter is replaced with a thinner one. The mount is also changed such that the transmitted beam can be now coupled to an other photodiode (earlier  the transmitted light was blocked by the mount). One more photodiode (PDA55) is introduced .So now the two photodiodes in the setup are PDA520 and PDA 55. 
 
 
ELOG V3.1.3-